OpenStudy (anonymous):

Can you pinpoint a mistake in my steps? Rewrite the following without logarithms. \[\ln I = \ln(2V) - [\ln(KR+r) - \ln K + KL]\] 1. \[2V \div [\ln(KR + r) - \ln K + KL]\] 2. \[ 2V \div [(KR + r) e^{KL} \div \ln K]\] 3. \I = 2VK \div (KR + r)e ^{KL} \

OpenStudy (anonymous):

What were the steps you took?

OpenStudy (anonymous):

\[\ln I = \ln(2V) - [\ln(KR+r) - \ln K + KL]\] \[2V \div [\ln(KR + r) - \ln(K) + KL]\] \[ 2V \div [(KR + r) e^{KL} \div \ln K]\] \[I = 2VK \div (KR + r)e ^{KL}\]

OpenStudy (anonymous):

First line is the equation. The 3 other lines are my steps.

myininaya (myininaya):

\[lnI=\ln(2V)-\ln(K(KR+r))-KL=\ln (\frac{2V}{K(KR+r)})-KL\]

myininaya (myininaya):

\[lnI=\ln(\frac{2V}{K(KR+r)})-KL\]

OpenStudy (anonymous):

Raising e to the power of both sides from there, we get: \[I = e^{\ln(\frac{2V}{K(KR+r)})-KL}\] \(=e^{\ln(\frac{2V}{K(KR+r)})}e^{-KL}\) \(=\frac{2V}{K(KR+r)}e^{-KL}\)

myininaya (myininaya):

\[e^{lnI}=e^{\ln(\frac{2V}{K(KR+r)})-KL}\] \[I=e^{\ln (\frac{2V}{K(KR+r)})}e^{-KL}=\frac{2V}{K(KR+r)}e^{-KL}\]

OpenStudy (anonymous):

Oh, and we can distribute the denominator a bit better there to look nicer.. \[={2Ve^{-KL} \over K^2R + Kr}\]

myininaya (myininaya):

\[I=\frac{2V}{e^{KL}(K(KR+r)}\]

OpenStudy (anonymous):

Oh, yeah.. Better still.

myininaya (myininaya):

i forgot a closing parenthesis at the end

OpenStudy (anonymous):

\[={2V \over e^{KL}(K^2R + Kr)}\]

OpenStudy (anonymous):

Myin, you know you can right click someones expression and view the source to copy/paste right?

OpenStudy (anonymous):

Saves having to retype long complicated messes like that when you've already done it once.

myininaya (myininaya):

i didn't know thanks

OpenStudy (anonymous):

Thanks guys!